OpenStudy (anonymous):

Two triangular roofs are similar. The ratio of the corresponding sides of these roofs is 2:3. If the altitude of the smaller roof is 4.5 feet, find the corresponding altitude of the larger roof.

OpenStudy (anonymous):

4.5 ( 3/2)= 13.5/2= 6.75